# Optimal Error Analysis of Channel Estimation for IRS-assisted MIMO Systems

### v1

@@ Abstract
As intelligent reflecting surface (IRS) has emerged as a new and promising technology capable of configuring the wireless environment favorably, channel estimation for IRS-assisted multiple-input multiple-output (MIMO) systems has garnered extensive attention in recent years. While various algorithms have been proposed to address this challenge, there is a lack of rigorous theoretical error analysis. This paper aims to address this gap by providing theoretical guarantees in terms of stable recovery of channel matrices for noisy measurements. We begin by establishing the equivalence between IRS-assisted MIMO systems and a compact tensor train (TT)-based tensor-on-tensor (ToT) regression. Building on this equivalence, we then investigate the restricted isometry property (RIP) for complex-valued subgaussian measurements. Our analysis reveals that successful recovery hinges on the relationship between the number of user terminals (in the uplink scenario) or base stations (in the downlink scenario) and the number of time slots during which channel matrices remain invariant. Utilizing the RIP condition, we analyze the theoretical recovery error for the solution to a constrained least-squares optimization problem, including upper error bound and minimax lower bound, demonstrating that the error decreases inversely with the number of time slots and increases proportionally with the number of unknown elements in the channel matrices. In addition, we extend our error analysis to two more specialized IRS-assisted MIMO systems, incorporating low-rank channel matrices or an unknown IRS. Furthermore, we explore a multi-hop IRS scheme and analyze the corresponding recovery errors. Finally, we introduce and implement two nonconvex optimization algorithms–alternating least squares and alternating gradient descent–to validate our conclusions through simulations.
Index Terms:
Intelligent reflecting surface, restricted isometry property, error analysis, tensor train-based tensor-on-tensor regression.
@@ I Introduction
Considering the wireless environment as an additional variable for optimization holds the promise of significant performance gains. This promise is realized through the innovative concept of intelligent reflecting surface (IRS) [2, 3, 4, 5, 6, 7, 8, 9], which can dynamically reshape the wireless propagation environment to exhibit more desirable characteristics. An IRS is essentially a 2D surface composed of numerous tunable units, which can be implemented using cost-effective antennas or metamaterials and controlled in real-time to manipulate communication channels without emitting signals of their own. Recently, IRS-assisted communications have garnered significant attention [10], owing to their potential to enhance efficiency, communication range, and capacity in wireless communication systems. Nevertheless, to fulfill the role of an IRS and achieve efficient and dependable wireless communication, the acquisition of precise channel state information using channel estimation techniques is imperative, posing a formidable challenge. One particular hurdle stems from the assumption that an IRS typically comprises passive elements, thereby necessitating the receiver to estimate the cascaded channel based on pilots transmitted by the transmitter through the IRS. Recently, numerous matrix-based methods for channel estimation [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] have been proposed and demonstrated commendable performance in practical implementation.
@@
Although various optimization algorithms for IRS-assisted MIMO systems have demonstrated excellent performance, existing analyses of stable recovery focus solely on the required number of time slots by exploiting the uniqueness of solutions in matrix-based models [36] or the uniqueness conditions of tensor decompositions [30]. Despite these advances, there is no comprehensive error analysis for channel estimation in IRS-assisted MIMO systems, leading to the following critical question:
@@
Question: What is the optimal recovery error for the channel estimation of the IRS-assisted MIMO system?
Our contribution:
In this paper, we affirmatively address the main question through error analyses in the tensor-based muliple-access/multi-user IRS-assisted MIMO system. Instead of using CP or Tucker decompositions, we first rigorously establish that the combination of channel matrices and IRS is equivalent to the tensor train (TT) format [37], which has been extensively utilized in quantum physics for efficiently representing quantum states [38, 39, 40]. Additionally, we demonstrate that the IRS-assisted MIMO model aligns with the tensor-on-tensor (ToT) regression model [41, 42, 43] without any assumptions regarding the pilot signals. This mode extends the classical matrix/tensor regression framework, in which the response is the vector [44, 45, 46, 47, 48, 49], to accommodate transmitted signals (represented by the second word “tensor” in the name) and the received signals (represented by the second word “tensor” in the name) as tensors. Building on the definitions outlined above, the channel estimation in a standard IRS-assisted MIMO model can be equivalently viewed as a recovery problem within the TT-based ToT regression model. Importantly, while other matrix/tensor-based models can be equivalently transformed into this framework, the TT-based ToT regression model is particularly advantageous because its canonical TT format [50, Theorem 1] facilitates theoretical analysis and the ToT model provides a compact representation that streamlines the design of optimization methods tailored to various pilot signal configurations. Consequently, this paper primarily focuses on analyzing this model.
@@
Building on the previous discussion, our problem is reformulated as a recovery problem within the TT-based ToT regression model, representing a multiple-access/multi-user MIMO communication system assisted by an IRS in an uplink communication scenario. Towards that goal, we initially establish the restricted isometry property (RIP) for complex-valued subgaussian measurements, where each pilot signal element is a complex-valued subgaussian random variable. This analysis reveals an optimal size relationship between user terminals (UTs) and time slots. Notably, for a downlink communication scenario, RIP relates to base stations (BSs) and time slots. Subsequently, we investigate channel matrices recovery of the least square optimization problem from subgaussian measurements and derive recovery bounds relative to the number of time slots. Our findings illustrate that recovery error is inversely proportionate to the number of time slots, and successful recovery depends on the relationship between UTs and time slots. Aside from aforementioned constraints, our recovery analyses remain free from other coefficient constraints. Moreover, we extend recovery error considerations to include low-rank channel matrices [30] and unknown IRS [51]. Next, we explore a multi-hop IRSs scheme and analyze corresponding recovery errors, laying the groundwork for future multi-hop IRSs communication design. Finally, we propose two nonconvex optimization algorithms–alternating least squares and alternating gradient descent and simulations affirm the validity of theoretical analyses.
Notations
@@
We use calligraphic letters (e.g., $\mathcal{Y}$) to denote tensors, bold capital letters (e.g., $\bm{Y}$) to denote matrices, except for $\bm{X}_{i}$ which denotes the $i$-th order-$3$ tensor factors in the TT format ($i=2,\dots,N+M-1$), bold lowercase letters (e.g., $\bm{y}$) to denote vectors, and italic letters (e.g., $y$) to denote scalar quantities. $a^{*}$ is the complex conjugate of $a$. Elements of matrices and tensors are denoted in parentheses, as in Matlab notation. For example, $\mathcal{X}(s_{1},s_{2},s_{3})$ denotes the element in position $(s_{1},s_{2},s_{3})$ of the order-3 tensor $\mathcal{X}$. The inner product of $\mathcal{A}\in\mathbb{R}^{d_{1}\times\dots\times d_{N}}$ and $\mathcal{B}\in\mathbb{R}^{d_{1}\times\dots\times d_{N}}$ can be denoted as $\langle\mathcal{A},\mathcal{B}\rangle=\sum_{s_{1}=1}^{d_{1}}\cdots\sum_{s_{N}=1}^{d_{N}}\mathcal{A}(s_{1},\dots,s_{N})\mathcal{B}(s_{1},\dots,s_{N})$. $\|\mathcal{X}\|_{F}=\sqrt{\langle\mathcal{X},\mathcal{X}\rangle}$ is the Frobenius norm of $\mathcal{X}$. Tensor contraction is defined as $\mathcal{A}\times_{i}^{j}\mathcal{B}$ of size $d_{1}\times\cdots\times d_{i-1}\times\times d_{i+1}\times\cdots\times d_{N}\times h_{1}\times\cdots\times h_{j-1}\times h_{j+1}\times\cdots\times h_{M}$ with $(s_{1},\dots,s_{i-1},s_{i+1},\dots,s_{N},f_{1},\dots,f_{j-1},f_{j+1},\dots,f_{M})$-th entry being $\sum_{k}\mathcal{A}(s_{1},\dots,s_{i-1},k,s_{i+1},\dots,s_{N})\mathcal{B}(f_{1},\dots,\\ f_{j-1},k,f_{j+1},\dots,f_{M})$ for $\mathcal{A}\in\mathbb{R}^{d_{1}\times\cdots\times d_{i-1}\times d_{k}\times d_{i+1}\times\cdots\times d_{N}}$ and $\mathcal{B}\in\mathbb{R}^{h_{1}\times\cdots\times h_{j-1}\times d_{k}\times h_{j+1}\times\cdots\times h_{M}}$. The procedure of $\mathcal{A}\times_{i_{1},\dots,i_{n}}^{j_{1},\dots,j_{n}}\mathcal{B}$ can be viewed as a sequence of $n$ tensor contractions $\mathcal{A}\times_{i_{k}}^{j_{k}}\mathcal{B},k\in[n]$. $\|\bm{X}\|$ represents the spectral norm of the matrix $\bm{X}$ For a positive integer $K$, $[K]$ denotes the set $\{1,\dots,K\}$. For two positive quantities $a,b\in\mathbb{R}$, the inequality $b=O(a)$ means $b\leq ca$ for some universal constant $c$; likewise, $b=\Omega(a)$ indicates that $b\geq ca$ for some universal constant $c$.
@@ II Single-Hop IRS-Assisted MIMO System
Model formulation
We consider a muliple-access/multi-user MIMO communication system assisted by an IRS, in which the $P$ base stations (BSs) receive the signals transmitted by $U$ user terminals (UTs) via the IRS. The terminology employed in this paper is based on an uplink communication scenario, where the transmitter is referred to as the UT and the receiver as the BS. However, it is important to note that our signal model and theoretical analysis can be equally applied to the downlink case by merely reversing the roles of the transmitter and the receiver. Without loss of generality, we assume that each BS and each UT are equipped with the same number of antennas, denoted as $L$ and $M$ respectively. The IRS is composed of $N$ passive elements, or unit cells, capable of individually adjusting their reflection coefficients (i.e., phase shifts). Due to unfavorable propagation conditions, BS-UT channels are neglected [29]. We also assume quasi-static flat fading channel model, where all the channels remain invariant during $T$ time slots. Assuming a block-fading channel, the received signal model is usually given as follows [15, 29, 30, 31]
@@
The channel coherence time $T_{s}$ is divided into $K$ blocks, where each block has $T$ time slots so that $T_{s}=KT$ [29]. Let us define $\bm{y}(k,t)=\bm{y}((k-1)T+t)$ as the received signal at the $t$-th time slot of the $k$-th block, where $t\in[T]$ and $k\in[K]$. In addition, we denote $\bm{x}(k,t)$ and $\bm{s}(k,t)$ as the pilot signal and phase shift vectors associated with the $t$-th time slot of the $k$-th block. In each block, the user transmits the same pilot signal and the IRS phase shift vector is invariant, but the phase shift vector between blocks is changeable, which means that the IRS has $K$ phase configurations. Consequently, we have $\bm{s}(k)=\bm{s}(k,t)$ and $\bm{x}(t)=\bm{x}(k,t)$. The received signal is given by
@@
Compact tensor representation
To accommodate a more general setting for the pilot signal matrix $\bm{X}$, we can reinterpret (2) as a tensor-on-tensor (ToT) regression model. Specifically, by aggregating the received data over $K$ blocks, we can define three tensors as $\mathcal{Y}\in\mathbb{C}^{LP\times K\times T}$, $\mathcal{S}\in\mathbb{C}^{N\times K\times N}$ and $\mathcal{W}\in\mathbb{C}^{LP\times K\times T}$, where $\mathcal{Y}(:,k,:)=\bm{Y}(k)$, $\mathcal{S}(:,k,:)=\text{diag}(\bm{s}(k))$ and $\mathcal{W}(:,k,:)=\bm{W}(k)$. Then, the ground-truth channel tensor $\mathcal{B}^{\star}\in\mathbb{C}^{LP\times K\times UM}$ can be defined as follows:
[…]
@@
Notice that, unlike the construction of CP decomposition, which requires transforming (3) as demonstrated in [29, 30, 52, 13, 31], (3) is inherently represented in the standard tensor train (TT) format as each element of $\mathcal{B}^{\star}$ can be expressed in a matrix product form (a detailed introduction to the TT decomposition is provided in Appendix A). By applying the tensor contraction operation [53], the TT form of $\mathcal{B}^{\star}$ can be expressed as:
@@
[…]
where (×$(\mathcal{B}^{\star}\times_{3}^{1}\bm{X})(p,k,t)=\sum_{j=1}^{UM}\mathcal{B}^{\star}(p,k,j)\bm{X}(j,t)$. The linear map $\mathcal{X}(\mathcal{B}^{\star}):\mathbb{C}^{LP\times K\times UM}\rightarrow\mathbb{C}^{LP\times K\times T}$ models the measurement process.
@@
^$\widehat{\mathcal{B}}=\operatorname*{\text{arg~{}min}}_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\frac{1}{T}\|\mathcal{X}(\mathcal{B})-\mathcal{Y}\|_{F}^{2},$ (7)
@@
[…]
@@
The advantage of the least-squares formulation is that it allows the application of well-established tools from standard matrix/tensor regression models, such as the Restricted Isometry Property (RIP) condition [54, 55, 56, 57, 40], as detailed in the following discussion. Subsequently, we provide a rigorous theoretical analysis, deriving error bounds for channel estimation.
Error analysis
One advantage of viewing the channel estimation as the regression model is that analogous to the analysis in standard matrix/tensor regression models, a larger number of time slots $T$ tends to enhance denoising effectiveness since the error bound of channel estimation in the IRS-assisted MIMO model should heavily rely on $T$. Recently, we analyzed the statistical guarantee of the TT-based ToT regression model in [58]. However, these results cannot be directly applied to the IRS-assisted MIMO model because recovery errors pertain to the entire tensor recovery. Since the phase shift tensor $\mathcal{S}$ is often known, only the channel matrices are affected by noise. Consequently, the recovery error should be primarily related to the channel matrices. Next, we will establish the upper bound and the minimax lower bound of the recovery error for channel estimation.
To facilitate the recovery of channels $\bm{H}$ and $\bm{G}$ from their linear measurements $\mathcal{Y}$, the pilot signal (regression/measurement operator) need adhere to specific properties. One such desirable property is the RIP, extensively explored and popularized in compressive regression literature [54, 55, 56, 59]. This property has also been extended to structured tensors [57, 40].

### v2

@@ Abstract
As intelligent reflecting surface (IRS) has emerged as a new and promising technology capable of configuring the wireless environment favorably, channel estimation for IRS-assisted multiple-input multiple-output (MIMO) systems has garnered extensive attention in recent years. Despite the development of numerous algorithms to address this challenge, a comprehensive theoretical characterization of the optimal recovery error is still lacking. This paper aims to address this gap by providing theoretical guarantees in terms of stable recovery of channel matrices for noisy measurements. We begin by establishing the equivalence between IRS-assisted MIMO systems in the uplink scenario and a compact tensor train (TT)-based tensor-on-tensor (ToT) regression. Building on this equivalence, we then investigate the restricted isometry property (RIP) for complex-valued subgaussian measurements. Our analysis reveals that successful recovery hinges on the relationship between the number of user terminals and the number of time slots during which channel matrices remain invariant. Utilizing the RIP condition, we establish a theoretical upper bound on the recovery error for solutions to the constrained least-squares optimization problem, as well as a minimax lower bound for the considered model. Our analysis demonstrates that the recovery error decreases inversely with the number of time slots, and increases proportionally with the total number of unknown entries in the channel matrices, thereby quantifying the fundamental trade-offs in channel estimation accuracy. In addition, we explore a multi-hop IRS scheme and analyze the corresponding recovery errors. Finally, we have performed numerical experiments to support our theoretical findings.
@@ I Introduction
Considering the wireless environment as an additional variable for optimization holds the promise of significant performance gains. This promise is realized through the innovative concept of intelligent reflecting surface (IRS) [2, 3, 4, 5, 6, 7, 8, 9], which can dynamically reshape the wireless propagation environment to exhibit more desirable characteristics. An IRS is essentially a 2D surface composed of numerous tunable units, which can be implemented using cost-effective antennas or metamaterials and controlled in real-time to manipulate communication channels without emitting signals of their own. Recently, IRS-assisted communications have garnered significant attention [10], owing to their potential to enhance efficiency, communication range, and capacity in wireless communication systems. Nevertheless, to fulfill the role of an IRS and achieve efficient and dependable wireless communication, the acquisition of precise channel state information using channel estimation techniques is imperative, posing a formidable challenge. One particular hurdle stems from the assumption that an IRS typically comprises passive elements, thereby necessitating the receiver to estimate the cascaded channel based on pilots transmitted by the transmitter through the IRS. Recently, numerous matrix-based methods for channel estimation [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] have been proposed and demonstrated commendable performance in practical implementation.
@@
Various optimization algorithms for IRS-assisted MIMO systems have exhibited strong empirical performance. However, the theoretical analysis of stable recovery has largely focused on identifying the required number of time slots by leveraging conditions that ensure the uniqueness of the solution [31, 37, 33]. A lower bound on the estimation error based on the Cramér–Rao bound (CRB) was established in [31]. However, the lower bound could be suboptimal since it is obtained by analyzing a convex surrogate of the original nonconvex least-squares problem, which alters the intrinsic structure of the estimation problem and, consequently, fails to capture its underlying degrees of freedom (see Section II-D for a detailed discussion). To the best of our knowledge, a rigorous optimal error analysis for channel estimation in IRS-assisted MIMO systems is still lacking. This raises the following fundamental question:
@@
Question: What is the optimal recovery error for channel estimation in IRS-assisted MIMO systems, and can this limit be achieved by appropriately formulating the underlying estimation problem?
I-A Our contribution:
@@
In this paper, we affirmatively address the main question through error analyses in the muliple-access/multi-user IRS-assisted MIMO system. We begin by formally establishing that the combination of the channel matrices and the IRS structure can be naturally represented in the tensor train (TT) format [38]. Based on this equivalence, we reformulate the IRS-assisted MIMO model as a TT-based tensor-on-tensor (ToT) regression model [39, 40]. This reformulation simplifies our analysis and naturally leads to a recovery problem under the TT-based ToT regression model, which captures the uplink communication scenario in a multiple-access/multi-user IRS-assisted MIMO system. To advance this analysis, we initially establish the restricted isometry property (RIP) for complex-valued subgaussian measurements, where each pilot signal element is a complex-valued subgaussian random variable. This analysis reveals an optimal size relationship between user terminals (UTs) and time slots. We then investigate the recovery of channel matrices via constrained least squares minimization from subgaussian measurements, deriving optimal recovery error bounds that are proportional to the degrees of freedom of the underlying channels. Specifically, we establish an upper bound on the estimation error for constrained least squares objective and further derive a minimax lower bound that depends solely on the underlying IRS-assisted MIMO model, without relying on any specific loss function, thereby further highlighting the optimality of the upper bound. Compared to the CRB, the minimax lower bound more accurately reflects the intrinsic nonconvexity of the model, yielding a tight characterization of the recovery error. Moreover, we extend recovery error considerations to include low-rank channel matrices [31], sparse matrices [41] and unknown IRS [42]. Next, we explore a multi-hop IRSs scheme and analyze corresponding recovery errors, laying the groundwork for future multi-hop IRSs communication design. Finally, simulations affirm the validity of theoretical analyses.
I-B Notations
@@
We use calligraphic letters (e.g., $\mathcal{Y}$) to denote tensors, bold capital letters (e.g., $\bm{Y}$) to denote matrices, except for $\bm{X}_{i}$ which denotes the $i$-th order-33 tensor factors in the TT format ($i=2,\dots,N+M-1$), bold lowercase letters (e.g., $\bm{y}$) to denote vectors, and italic letters (e.g., $y$) to denote scalar quantities. $a^{*}$ is the complex conjugate of $a$. Elements of matrices and tensors are denoted in parentheses, as in Matlab notation. For example, $\mathcal{X}(s_{1},s_{2},s_{3})$ denotes the element in position $(s_{1},s_{2},s_{3})$ of the order-3 tensor $\mathcal{X}$. The inner product of $\mathcal{A}\in\mathbb{R}^{d_{1}\times\dots\times d_{N}}$ and $\mathcal{B}\in\mathbb{R}^{d_{1}\times\dots\times d_{N}}$ can be denoted as $\langle\mathcal{A},\mathcal{B}\rangle=\sum_{s_{1}=1}^{d_{1}}\cdots\sum_{s_{N}=1}^{d_{N}}\mathcal{A}(s_{1},\dots,s_{N})\mathcal{B}(s_{1},\dots,s_{N})$. $\|\mathcal{X}\|_{F}=\sqrt{\langle\mathcal{X},\mathcal{X}\rangle}$ is the Frobenius norm of $\mathcal{X}$. Tensor contraction is defined as $\mathcal{A}\times_{i}^{j}\mathcal{B}$ of size $d_{1}\times\cdots\times d_{i-1}\times\times d_{i+1}\times\cdots\times d_{N}\times h_{1}\times\cdots\times h_{j-1}\times h_{j+1}\times\cdots\times h_{M}$ with $(s_{1},\dots,s_{i-1},s_{i+1},\dots,s_{N},f_{1},\dots,f_{j-1},f_{j+1},\dots,f_{M})$-th entry being $\sum_{k}\mathcal{A}(s_{1},\dots,s_{i-1},k,s_{i+1},\dots,s_{N})\mathcal{B}(f_{1},\dots,\\ f_{j-1},k,f_{j+1},\dots,f_{M})$ for $\mathcal{A}\in\mathbb{R}^{d_{1}\times\cdots\times d_{i-1}\times d_{k}\times d_{i+1}\times\cdots\times d_{N}}$ and $\mathcal{B}\in\mathbb{R}^{h_{1}\times\cdots\times h_{j-1}\times d_{k}\times h_{j+1}\times\cdots\times h_{M}}$. The procedure of $\mathcal{A}\times_{i_{1},\dots,i_{n}}^{j_{1},\dots,j_{n}}\mathcal{B}$ can be viewed as a sequence of $n$ tensor contractions $\mathcal{A}\times_{i_{k}}^{j_{k}}\mathcal{B},k\in[n]$. The symbol * denotes the Khatri-Rao product, and the operator $\operatorname*{vec}(\cdot)$ refers to the vectorization of a matrix. $\|\bm{X}\|$ and $\|\bm{X}\|_{0}$ respectively represent the spectral norm and $\ell_{0}$ norm of the matrix $\bm{X}$. For a positive integer $K$, $[K]$ denotes the set $\{1,\dots,K\}$. For two positive quantities $a,b\in\mathbb{R}$, the inequality $b=O(a)$ means $b\leq ca$ for some universal constant $c$; likewise, $b=\Omega(a)$ indicates that $b\geq ca$ for some universal constant $c$.
@@ II Single-Hop IRS-Assisted MIMO System
II-A Model formulation
We consider a muliple-access/multi-user MIMO communication system assisted by an IRS, in which the $P$ base stations (BSs) receive the signals transmitted by $U$ UTs via the IRS. The terminology employed in this paper is based on an uplink communication scenario, where the transmitter is referred to as the UT and the receiver as the BS. Without loss of generality, we assume that each BS and each UT are equipped with the same number of antennas, denoted as $L$ and $M$ respectively. The IRS is composed of $N$ passive elements, or unit cells, capable of individually adjusting their reflection coefficients (i.e., phase shifts). Due to unfavorable propagation conditions, BS-UT channels are neglected [30]. We also assume quasi-static flat fading channel model, where all the channels remain invariant during $T$ time slots. Assuming a block-fading channel, the received signal model is usually given as follows [15, 30, 31, 32]
@@ II-A Model formulation
The channel training time $T_{s}$ is divided into $K$ blocks, where each block has $T$ time slots so that $T_{s}=KT$ [30]. Let us define $\bm{y}(k,t)=\bm{y}((k-1)T+t)$ as the received signal at the $t$-th time slot of the $k$-th block, where $t\in[T]$ and $k\in[K]$. Similarly, we denote $\bm{x}(k,t)$, $\bm{s}(k,t)$ and $\bm{w}(k,t)$ as the pilot signal, phase shift and noise vectors associated with the $t$-th time slot of the $k$-th block. We adopt the structured time-domain training protocol proposed in [30], wherein the IRS phase shift vector remains fixed throughout the TT time slots within each block and changes from block to block. Meanwhile, the pilot signals $\begin{bmatrix}\bm{x}(1)\cdots\bm{x}(T)\end{bmatrix}$ are identically repeated across all $K$ blocks. Consequently, we have
$\bm{s}(k)$ $\!\!\!\!=\!\!\!\!$ $\bm{s}(k,t),t\in[T],$
$\bm{x}(t)$ $\!\!\!\!=\!\!\!\!$ $\bm{x}(k,t),k\in[K].$
Under these assumptions, the received signal in (1) is given by
@@
II-B Compact tensor representation
To accommodate a more general setting for the pilot signal matrix $\bm{X}$ and facilitate the analysis of the optimal recovery error in channel estimation, we can reinterpret (2) as a tensor-on-tensor (ToT) regression model [44, 45, 39, 40] which generalizes the classical matrix/tensor regression framework where the response variable is a vector [46, 47, 48, 49, 50, 51, 52, 53]. Specifically, by aggregating the received data over $K$ blocks, we can define three tensors as $\mathcal{Y}\in\mathbb{C}^{LP\times K\times T}$, $\mathcal{S}\in\mathbb{C}^{N\times K\times N}$ and $\mathcal{W}\in\mathbb{C}^{LP\times K\times T}$, where $\mathcal{Y}(:,k,:)=\bm{Y}(k)$, $\mathcal{S}(:,k,:)=\text{diag}(\bm{s}(k))$ and $\mathcal{W}(:,k,:)=\bm{W}(k)$. Then, the ground-truth channel tensor $\mathcal{B}^{\star}\in\mathbb{C}^{LP\times K\times UM}$ can be defined as follows:
$\mathcal{B}^{\star}(p,k,m)$ $\!\!\!\!=\!\!\!\!$ $\bm{H}(p,:)\mathcal{S}(:,k,:)\bm{G}(:,m)$ (3)
$\!\!\!\!=\!\!\!\!$ $\bm{H}(p,:)\text{diag}(\bm{s}(k))\bm{G}(:,m).$
@@ II-B Compact tensor representation
Notice that, unlike the construction of CP decomposition, which requires transforming (3) as demonstrated in [30, 31, 43, 13, 32], (3) is inherently represented in the standard tensor train (TT) format [38] as each element of $\mathcal{B}^{\star}$ can be expressed in a matrix product form (a detailed introduction to the TT decomposition is provided in the Appendix A). By applying the tensor contraction operation [54], the TT form of $\mathcal{B}^{\star}$ can be expressed as:
@@
$\mathcal{Y}=\mathcal{X}(\mathcal{B}^{\star})+\mathcal{W}$ $\!\!\!\!=\!\!\!\!$ $\mathcal{B}^{\star}\times_{3}^{1}\bm{X}+\mathcal{W}$ (6)
$\!\!\!\!=\!\!\!\!$ $[\bm{H},\mathcal{S},\bm{G}]\times_{3}^{1}\bm{X}+\mathcal{W},$
where $(\mathcal{B}^{\star}\times_{3}^{1}\bm{X})(p,k,t)=\sum_{j=1}^{UM}\mathcal{B}^{\star}(p,k,j)\bm{X}(j,t)$. The linear map $\mathcal{X}(\mathcal{B}^{\star}):\mathbb{C}^{LP\times K\times UM}\rightarrow\mathbb{C}^{LP\times K\times T}$ models the measurement process. Here, we emphasize that the TT-based ToT regression model in (6) serves solely to simplify the theoretical analysis and preserves all the information contained in the original model (2).
@@
$\widehat{\mathcal{B}}=\operatorname*{\text{arg~min}}_{\mathcal{B}\in\mathbb{B}_{N,\mathcal{S}}}\frac{1}{T}\|\mathcal{X}(\mathcal{B})-\mathcal{Y}\|_{F}^{2},$ (7)
@@
$\mathbb{B}_{N,\mathcal{S}}$ $\!\!\!\!=\!\!\!\!$ $\{\mathcal{B}=[\bm{H},\mathcal{S},\bm{G}]:\bm{H}\in\mathbb{C}^{LP\times N},\bm{G}\in\mathbb{C}^{N\times UM}\$(8)
@@
The advantage of the least-squares formulation is that it allows the application of well-established tools from standard matrix/tensor regression models, such as the RIP condition [55, 56, 57, 58, 50], as detailed in the following discussion. Subsequently, we provide a rigorous theoretical analysis, deriving error bounds for channel estimation.
II-C Optimal error analysis
One advantage of viewing the channel estimation as the regression model is that analogous to the analysis in standard regression models, a larger number of time slots $T$ tends to enhance denoising effectiveness since the error bound of channel estimation in the IRS-assisted MIMO model should heavily rely on $T$. Recently, we analyzed the statistical guarantee of the TT-based ToT regression model in [40]. Compared with that model, the IRS-assisted MIMO setting introduces additional structural constraints, as the phase-shift tensor $\mathcal{S}$ is typically known and only the channel matrices are corrupted by noise. Consequently, a direct extension of the results in [40] would lead to non-tight recovery bounds that do not properly scale with the degrees of freedom in the unknown channel matrices. Since the recovery error is primarily determined by the channel matrices, we next establish both the upper and minimax lower bounds for the corresponding channel estimation problem.
To facilitate the recovery of channels $\bm{H}$ and $\bm{G}$ from their linear measurements $\mathcal{Y}$, the pilot signal (regression/measurement operator) need adhere to specific properties. One such desirable property is the RIP, extensively explored and popularized in compressive regression literature [55, 56, 57, 59], and later extended to structured tensor models [58, 50]. From a theoretical perspective, the RIP is typically established by assuming that the measurement operators are drawn randomly from a suitable distribution, with subgaussian measurement ensembles serving as a canonical example.